OpenStudy (anonymous):

8th grade math help? Fan and medal 5 questions. :)

OpenStudy (anonymous):

Find the slope of the following graphs https://www.connexus.com/content/media/680582-532012-11420-PM-2043777340.png a. 1/2 b. 2 c. -2 d. - 1/2

OpenStudy (love_ranaa):

Access Denied.. You have to login to view the question && I don't have a login :)

OpenStudy (yolo_king):

Could you take a screen shot ?

OpenStudy (anonymous):

I saved it to my computer and attached it as a file.

OpenStudy (hulagirlmb):

Okay, you see the points? Pick 2

OpenStudy (hulagirlmb):

Now, do you know the equation for slope?

OpenStudy (anonymous):

/.\ Idk :c If I was told how to do this, I wouldnt have any problem

OpenStudy (anonymous):

The formula for slope is:\[\frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (hulagirlmb):

Yeah haha. Couldn't figure out how to do that

OpenStudy (anonymous):

Choose two sets of coordinates \[(x_1, y_1) (x_2, y_2)\]

OpenStudy (anonymous):

The set I'm choosing looks like : (2, -2) & (1,0). The equation is \[\frac{ 0 + 2 }{ 1 - 2 }\]

OpenStudy (anonymous):

After solving this, you see that \[\frac{ 2 }{ -1 }\] is what is left.

OpenStudy (anonymous):

Seeing this, you know the answer is -2.